# Regurgitative Training: The Value of Real Data in Training Large Language Models

### v1

@@ 3.2 Experiments with Models Trained from Scratch
We follow Vaswani et al., (2017) to build the baseline translation models. Transformer has an encoder-decoder architecture, which uses stacked layers of multi-head self-attention and point-wise, fully connected feed-forward networks for both encoder and decoder. It also employs a residual connection on each sub-layer, followed by layer normalization. For implementational details of these transformer elements, we refer to Vaswani et al., (2017).
As our previous fine-tuning results have shown, the performance impact of regurgitative training can vary with the capability of the model used to generate training data. Therefore, we train both a “low-performance” and a ”high-performance” baseline models. This is done by gradually adding 50,000 sentence pairs (randomly sampled from the German-English corpus) per batch for training, and evaluate the model’s translation performance on a fixed testing dataset of 50,000 sentence pairs. As shown in Figure 3, we observe that the model’s performance improves quickly with the initial increase in training data size, and saturates after being trained with sufficient data. We choose the model trained with 50,000 data instances as our low-performance baseline and the one trained with 500,000 data instances as the high-performance baseline.
@@
These two baseline models, corresponding to different performance levels, are then used to evaluate the effects of regurgitative training. We randomly sample a total of 300,000 data instances (outside of the training data of both the low- and high-performance baseline models) designated for regurgitative training. In batches of 10,000 data instances, we continue training both the low-performance baseline model and the high-performance baseline model with (i) real human-generated data, (ii) data generated by the low-performance model, and (iii) data generated by the high-performance model. After each batch of training, we evaluate the all models’ performance on the same testing data of 50,000 instances. The results are presented in Figure 4.
@@
For both low-performance and high-performance baseline models, regurgitative training with data generated from the low-performance model clearly underperforms training with real data. The same is true for regurgitative training of high-performance model with data generated by itself, though the performance gap is fairly small. Curiously, regurgitative training of low-performance model with data generated by high-performance model actually outperforms training with real data for the first 19 batches (i.e., top two lines in the right plot). To understand whether this is a sustainable performance advantage, we sample more data to carry out another 20 batches of regurgitative training in this case. The results, shown in Figure 5, show that regurgitative training performance starts to plateau around 30 batches, and underperforms training with real data thereafter.
@@
In reality, it is plausible that a mixture of both model-generated data and real data are used for training. We therefore carry out another set of experiments to check how the proportion of model-generated data in the mix affect model performance. We simulate five scenarios, where the proportion of model-generated data is 100%, 75%, 50%, 25%, and 0% respectively, and the rest of each mixture consists of real data. Naturally, the scenarios with 100% and 0% of model-generated data are the same as training purely with model-generated data or purely with real data. For clarity, we focus on training each baseline model with data generated by itself (mixed with different proportions of real data). Other experiment settings are the same as before, and the results are shown in Figure 6.
@@
In the case of high-performance baseline (i.e., left plot), because the performance gap between regurgitative training and training with real data is relatively small to begin with, the pattern is obfuscated by local performance fluctuations. Nonetheless, having a higher proportion of real data still generally leads to better performance. In contrast, the pattern becomes much clearer in the case of low-performance baseline (i.e., right plot). We can see that even a small amount of model-generated data is enough to slow down learning. As a higher proportion of model-generated data is used, the model’s performance continues to deteriorate.
@@ 3.3 Replication: Question Answering Task
In addition to machine translation, we also conduct a replication study with another common generative language task – Question Answering (Q&A). We use the Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al.,, 2016), which is a widely used benchmarking dataset for developing and testing Q&A methods. SQuAD is a reading comprehension dataset composed of questions created by crowd-workers based on a collection of Wikipedia articles, with answers being segments of texts from the corresponding passages in the articles. The dataset includes 87,599 entries in the training dataset (constructed from 442 articles) and 10,570 entries in the testing dataset (constructed from 48 other articles).
@@
Instead of end-to-end training (as was done in the previous section), here we use a pre-trained multilingual BERT model (bert-base-cased, Devlin et al.,, 2018) to extract word embeddings, which are fed into a feedforward neural network model for Q&A. Doing so allows us to evaluate regurgitative training under yet another widely adopted strategy for training generative language models (i.e., leveraging pre-trained representation models). [^5] This strategy is often also referred to as “fine-tuning” a pre-trained model. We refrain from using this term here in order to avoid confusion with our fine-tuning experiments in Section 3.1, which are performed on top of existing LLMs (rather than a BERT-like representation model). We follow Rajpurkar et al., (2016) to evaluate Q&A performance with two metrics: Exact Match and F-1 score. Exact match measures the percentage of predicted answers that match the ground-truth answers exactly, and F-1 score measures the overlap between the predicted answers and the ground-truth answers by treating both predictions and ground-truths as bags of tokens. We calculate the average F-1 over all questions in the testing data.
@@
Following the same procedure as in the previous section, we use increasing amounts of real data to train baseline models in order to identify a low-performance model and a high-performance model (see Appendix A for detailed results). The low-performance model is trained on all entries from 40 articles and achieves 70.77% exact match rate and 80.34% average F-1 score, whereas the high-performance model is trained on all entries from 200 articles and achieves 78.63% exact match rate and 86.59% average F-1 score.
@@
Next we use these two baseline models for regurgitative training on entries from the remaining 242 articles (not used in training the two baseline models). In batches of 10 articles, we continue training both the low- and high-performance baselines with (i) real human-generated data, (ii) data generated by the low-performance model, and (iii) data generated by the high-performance model, for a total of 20 batches. After each batch of training, we evaluate the all models’ performance on the same testing dataset provided by SQuAD. The results are included in Figure 7, where the first row shows performance of regurgitative training the high-performance baseline model and the second row shows performance of regurgitative training the low-performance baseline model.
@@ 4 Understanding Performance Loss from Regurgitative Training
Why does regurgitative training hurt performance compared to training with real data? In this section, we offer some preliminary evidence into the underlying mechanisms. Using the translation task as an example, we focus on characterizing the differences between LLM-generated training data and real data, and discuss how these differences may impact the performance.
@@
Although BLEU is widely used to measure translation quality, it also has an important limitation that it does not explicitly account for the semantic meaning of words. A translation that uses different words than those in the ground-truth will have a low BLEU score even if it is semantically correct. In other words, having a lower BLEU score does not necessarily mean that the translation is more erroneous. In light of this, we construct two new measures, both aiming at quantifying the semantic differences of LLM-generated data vs. real data. We take each set of training data (generated by one of the LLMs or human) and perform several pre-processing steps, including (i) lower-casing, (ii) removing punctuation, (iii) removing stopwords, and (iv) lemmatization (i.e., reducing a word to its stem form). These pre-processing steps allow us to focus only on the substantive content of each translation. [^6] For example, two different translations “Tomorrow will be raining!” and “Tomorrow will rain.” will both become “tomorrow rain” after pre-processing, as it lower-cases both sentences, removes punctuation, removes stopwords “will” and “be”, and reduces “raining” to its stem “rain”. The first metric is computed as the average cosine similarity between the embeddings of LLM and ground-truth translations, where the embeddings are obtained from the Sentence Transformer model (Reimers and Gurevych,, 2019). After pre-processing, a smaller cosine similarity implies greater semantic discrepancies of LLM translations from the ground-truths, which is indicative of translation errors. The second metric counts the number of word tokens in a ground-truth translation that satisfy two conditions: (1) they do not show up in the corresponding LLM translation and (2) even their synonyms (retrieved based on WordNet, Miller,, 1995) do not show up in the LLM translation. These non-synonymous deviations likely represent words mistranslated by LLM. Results of these two metrics are reported in the second and third rows of Table 1. We see that the two sets of GPT-generated data have higher semantic similarities with ground-truths and lower non-synonym deviations than LLAMA2, again supporting the mechanism that translation errors are partially responsible for the performance reduction of regurgitative training.
@@
Beyond errors, we also test a different mechanism related to lexical diversity. Several recent work suggest that LLM-generated content appears to be more homogeneous than human-generated content (Doshi and Hauser,, 2023; Anderson et al.,, 2024; Zhou and Lee,, 2024). We suspect that regurgitative training with less diverse LLM-generated data may hinder the model’s ability to generalize and result in lower testing performance. We quantify lexical diversity with two metrics. The first is a straightforward count of the total number of unique word tokens in ground-truth or LLM translations. The second adopts the self-BLEU metric proposed by Zhu et al., (2018). Self-BLEU is the BLEU score of a given text against all other texts in a corpus. Because BLEU captures lexical similarity, self-BLEU accordingly reflects how similar a text is with the rest of the corpus (higher self-BLEU implies lower diversity). For LLM-generated translations, errors may artificially decrease self-BLEU without meaningfully increase lexical diversity. We therefore remove the previously mentioned non-synonymous deviations (as approximation of errors) from LLM translations. We then average self-BLEU over the 5,000 training data points. Results of both metrics are reported in Table 2. Ground-truth translations consistently use more unique tokens and have lower average self-BLEU than LLM translations. GPT translations use more unique tokens than LLAMA2 and the three LLMs have similar average self-BLEU.
@@ 5 Mitigating Performance Loss from Regurgitative Training
Details of each strategy and the corresponding evaluations on the translation task are discussed in the rest of this section. [^7] As will be discussed later, the first, quality-based mitigation strategy naturally also works for the Q&A task, which we will demonstrate as part of Section 5.1.2. However, it is worth noting up front that the goal of mitigation is not to completely close the gap from the performance of training with real data – this may not be realistic in the short term. Instead, the goal is to use LLM-generated synthetic data in a more careful manner to reduce performance loss. [^8] One might suggest the best strategy to reduce performance loss is not to use any synthetic data at all. However, as we discussed in Section 1, real human-generated data alone may not be sufficient to train next-generation LLMs.
@@ 5.1 Mitigation Strategy based on Quality Quantification
The labeled dataset is randomly split into 80% for training and 20% for testing. Each instance of the labeled dataset is structured as $input=(g_{1},g_{2},\ldots,g_{M},[SEP],e_{1},e_{2},\ldots,e_{N},[SEP]),label=BLEU$, where $g_{i}$ represents tokens in German sentences, $e_{j}$ represents tokens in English translations, and $[SEP]$ denotes the special separation token. Similar to an approach used in Chowdhury et al., (2021), we derive embedding of the entire input sequence from BERT (with the bert-base-multilingual-uncased pre-trained model), which are then used as input features to eight different supervised learning techniques for BLEU prediction. We train a separate BLEU prediction models for each of the three LLMs, and the testing performance of these BLEU prediction models are summarized in Appendix B. We find that the Bayesian Ridge technique exhibits relatively superior BLEU prediction performance (achieving lower MAE and MSE values).
@@
$\text{Translation Entropy}=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v=1}^{V}p(y_{t}^{v})\log{p(y_{t}^{v})})$ (3)
@@
Finally, we also apply the quality-based mitigation strategy on the Q&A task. For a given question, we analogously derive an entropy score for a model-generated answer to reflect the uncertainty of the probability distribution over all possible answers. For Q&A task, the transformer model generates a candidate answer by predicting the positions of a start token and an end token in a given passage that potentially contains the answer (Rajpurkar et al.,, 2016). The start / end tokens then jointly determine the answer text. Therefore, the probability score of a candidate answer is the product of the probabilities associated with the start and end tokens, and the entropy score can be computed as
Answer $\text{Answer Entropy}=-\sum_{a=1}^{A}p(y_{a}^{start})p(y_{a}^{end})\log\left\{p(y_{a}^{start})p(y_{a}^{end})\right\}$ (4)
where $a\in A$ is a candidate answer; $p(y_{a}^{start})$ and $p(y_{a}^{end})$ respectively denote the predicted probabilities of the start / end tokens that demarcate answer $a$. Same as before, a lower entropy corresponds to a more confident model-generated answer.
We then follow the same settings in Section 3.3 to conduct regurgitative training on both low-performance and high-performance baseline models, except that the model-generated data are added in the order of answer entropy (from low to high). At the article level, we prioritize articles based on the lowest entropy value among their constituent answers. The results are shown in Figure 10.
@@
We again observe some performance gains of regurgitative training with quality-ranked data, but the differences tend to be fairly small. To summarize, while the experiments in this section generally support the potential of the quality-based mitigation, we note that the performance gap from training with real data remains substantial. Put differently, the value of real data in training language models cannot be substituted by quality-aware regurgitative training.

### v2

@@ 3.2 Experiments with Models Trained from Scratch
We follow Vaswani et al., (2017) to build the baseline transformer models. Transformer has an encoder-decoder architecture, which uses stacked layers of multi-head self-attention and point-wise, fully connected feed-forward networks for both the encoder and decoder. It also employs a residual connection on each sub-layer, followed by layer normalization. For implementational details of these transformer elements, we refer to Vaswani et al., (2017).
As our previous fine-tuning results have shown, the performance impact of regurgitative training can vary with the capability of the model used to generate training data. Therefore, we train both a “low-performance” and a ”high-performance” baseline models. This is done by gradually adding 50,000 sentence pairs (randomly sampled from the German-English corpus) per batch for training, and evaluate the model’s translation performance on a fixed testing dataset of 50,000 sentence pairs. As shown in Figure 3, we observe that the model’s performance improves quickly with the initial increase in training data size, and saturates after being trained with sufficient data. We choose the model trained with 50,000 data instances (i.e., 1 batch) as our low-performance baseline and the one trained with 500,000 data instances (i.e., 10 batches) as the high-performance baseline.
@@
These two baseline models, corresponding to different performance levels, are then used to evaluate the effects of regurgitative training. We randomly sample a total of 300,000 data instances (outside of the training data of both the low- and high-performance baseline models) designated for regurgitative training. In batches of 10,000 data instances, we continue training both the low-performance baseline model and the high-performance baseline model with (i) real human-generated data, (ii) data generated by the low-performance model, and (iii) data generated by the high-performance model. After each batch of training, we evaluate all models’ performances on the same testing data of 50,000 instances. The results are presented in Figure 4.
@@
For both low-performance and high-performance baseline models, regurgitative training with data generated from the low-performance model clearly underperforms training with real data. The same is true for regurgitative training of high-performance model with data generated by itself, though the performance gap is fairly small. Curiously, regurgitative training of low-performance model with data generated by high-performance model actually outperforms training with real data for the first 19 batches (i.e., top two lines in the right plot). To understand whether this is a sustainable performance advantage, we sample more data to carry out another 20 batches of regurgitative training in this case. The results, as seen in Figure 5, show that regurgitative training performance starts to plateau around 30 batches, and underperforms training with real data thereafter.
@@
In reality, it is plausible that a mixture of both model-generated data and real data are used for training. We therefore carry out another set of experiments to check how the proportion of model-generated data in the mix affects model performance. We simulate five scenarios, where the proportion of model-generated data is 100%, 75%, 50%, 25%, and 0% respectively, and the rest of each mixture consists of real data. Naturally, the scenarios with 100% and 0% of model-generated data are the same as training purely with model-generated data or purely with real data. For clarity, we focus on training each baseline model with data generated by itself (mixed with different proportions of real data). Other experiment settings are the same as before, and the results are shown in Figure 6.
@@
In the case of high-performance baseline (i.e., left plot), because the performance gap between regurgitative training and training with real data is relatively small to begin with, the pattern is obfuscated by local performance fluctuations. Nonetheless, having a higher proportion of real data still generally leads to better performance. In contrast, the pattern becomes much clearer in the case of low-performance baseline (i.e., right plot). We can see that even a small amount of model-generated data is enough to deteriorate learning. As a higher proportion of model-generated data is used, the model’s performance continues to deteriorate.
Finally, we note that because the high-/low-performance baseline models are initially trained with different volumes of data, the data added to the two models during regurgitative training naturally amount to different proportions of their initial training data – specifically, each batch of 10,000 data corresponds to 20% of low-performance model’s training data and 2% of high-performance model’s training data. As a robustness check, we repeat regurgitative training of the two models by adding a fixed percentage of each model’s training data, i.e., 10% per batch. The results, reported in Appendix A, remain highly consistent with our main findings above.
@@ 3.3 Replication: Question Answering Task
In addition to machine translation, we also conduct a replication study with another common generative language task – Question Answering (Q&A). We use the Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al.,, 2016), which is a widely used benchmarking dataset for developing and testing Q&A methods. SQuAD is a reading comprehension dataset composed of questions created by crowd-workers based on a collection of Wikipedia articles, with answers being segments of texts from the corresponding passages in the articles. The dataset includes 87,599 entries in the training dataset (constructed from 442 articles) and 10,570 entries in the development dataset (constructed from 48 other articles) which we use as testing data.
@@
Instead of end-to-end training (as was done in the previous section), here we use a pre-trained BERT model (bert-base-cased, Devlin et al.,, 2018) to extract word embeddings, which are fed into a feedforward neural network model for Q&A. Doing so allows us to evaluate regurgitative training under yet another widely adopted strategy for training generative language models (i.e., leveraging pre-trained representation models). [^5] This strategy is often also referred to as “fine-tuning” a pre-trained model. We refrain from using this term here in order to avoid confusion with our fine-tuning experiments in Section 3.1, which are performed on top of existing LLMs (rather than a BERT-like representation model). We follow Rajpurkar et al., (2016) to evaluate Q&A performance with two metrics: Exact Match and F-1 score. Exact match measures the percentage of predicted answers that match the ground-truth answers exactly, and F-1 score measures the overlap between the predicted answers and the ground-truth answers by treating both predictions and ground-truths as bags of tokens. We calculate the average F-1 score over all questions in the testing data.
@@
Following the same procedure as in the previous section, we use increasing amounts of real data to train baseline models in order to identify a low-performance model and a high-performance model (see Appendix B for detailed results). The low-performance model is trained on all entries from 40 articles and achieves 70.68% exact match rate and 80.38% average F-1 score, whereas the high-performance model is trained on all entries from 200 articles and achieves 78.47% exact match rate and 86.54% average F-1 score.
@@
Next we use these two baseline models for regurgitative training on entries from the remaining 242 articles (not used in training for the two baseline models). In batches of 10 articles, we continue training both the low- and high-performance baselines with (i) real human-generated data, (ii) data generated by the low-performance model, and (iii) data generated by the high-performance model, for a total of 20 batches. After each batch of training, we evaluate all models’ performances on the same testing dataset provided by SQuAD. The results are included in Figure 7, where the first row shows performance of regurgitative training the high-performance baseline model and the second row shows performance of regurgitative training the low-performance baseline model.
@@ 4 Understanding Performance Loss from Regurgitative Training
Why does regurgitative training hurt performance compared to training with real data? In this section, we offer some preliminary evidence into the underlying mechanisms. Using the translation task as an example, we focus on characterizing the differences between LLM-generated training data and real data, and discuss how these differences may impact the performance. Keeping in mind that theoretical / mathematical understandings of LLMs (and complex deep neural networks in general) still remain largely out of reach, the following explorations are empirical in nature.
@@
Although BLEU is widely used to measure translation quality, it also has an important limitation that it does not explicitly account for the semantic meaning of words. A translation that uses different words than those in the ground-truths will have a low BLEU score even if it is semantically correct. In other words, having a lower BLEU score does not necessarily mean that the translation is more erroneous. In light of this, we construct two new measures, both aiming at quantifying the semantic differences of LLM-generated data vs. real data. We take each set of training data (generated by one of the LLMs or human) and perform several pre-processing steps, including (i) lower-casing, (ii) removing punctuation, (iii) removing stopwords, and (iv) lemmatization (i.e., reducing a word to its stem form). These pre-processing steps allow us to focus only on the substantive content of each translation. [^6] For example, two different translations “Tomorrow will be raining!” and “Tomorrow will rain.” will both become “tomorrow rain” after pre-processing, as it lower-cases both sentences, removes punctuation, removes stopwords “will” and “be”, and reduces “raining” to its stem “rain”. The first metric is computed as the average cosine similarity between the embeddings of LLM-generated and ground-truth translations, where the embeddings are obtained from the Sentence Transformer model (Reimers and Gurevych,, 2019). After pre-processing, a smaller cosine similarity implies greater semantic discrepancies of LLM translations from the ground-truths, which is indicative of translation errors. The second metric counts the number of word tokens in a ground-truth translation that satisfy two conditions: (1) they do not show up in the corresponding LLM translation and (2) even their synonyms (retrieved based on WordNet, Miller,, 1995) do not show up in the LLM translation. These non-synonymous deviations likely represent words mistranslated by LLM. Results of these two metrics are reported in the second and third rows of Table 1. We see that the two sets of GPT-generated data have higher semantic similarities with ground-truths and lower non-synonym deviations than LLAMA2, again supporting the mechanism that translation errors are partially responsible for the performance reduction of regurgitative training.
@@
Beyond errors, we also test a different mechanism related to lexical diversity. Several recent work suggest that LLM-generated content appears to be more homogeneous than human-generated content (Doshi and Hauser,, 2023; Anderson et al.,, 2024; Zhou and Lee,, 2024). We suspect that regurgitative training with less diverse LLM-generated data may hinder the model’s ability to generalize and result in lower testing performance. We quantify lexical diversity with two metrics. The first is a straightforward count of the total number of unique word tokens in ground-truth or LLM translations. The second adopts the self-BLEU metric proposed by Zhu et al., (2018). Self-BLEU is the BLEU score of a given text against all other texts in a corpus. Because BLEU captures lexical similarity, self-BLEU accordingly reflects how similar a text is with the rest of the corpus (higher self-BLEU implies lower diversity). For LLM-generated translations, errors may artificially decrease self-BLEU without meaningfully increase lexical diversity. We therefore remove the previously mentioned non-synonymous deviations (as approximation of errors) from LLM translations. We then average self-BLEU over the 5,000 training data points. Results of both metrics are reported in Table 2. Ground-truth translations consistently use more unique tokens and have significantly lower average self-BLEU than LLM translations ($p<0.001$). GPT translations use a bit more unique tokens than LLAMA2 and the three LLMs have similar average self-BLEU (their self-BLEU differences are not statistically significant, $p>0.05$).
@@ 5 Mitigating Performance Loss from Regurgitative Training
Details of each strategy and the corresponding evaluations on the translation task are discussed in the rest of this section. The first quality-based mitigation strategy is also naturally applicable on the Q&A task, which we will demonstrate as part of Section 5.1.2. However, it is worth noting up front that the goal of mitigation is not to completely close the gap from the performance of training with real data – this may not be realistic in the short term. Instead, the goal is to use LLM-generated synthetic data in a more careful manner to reduce performance loss. [^7] One might suggest the best strategy to reduce performance loss is not to use any synthetic data at all. However, as we discussed in Section 1, real human-generated data alone may not be sufficient to train next-generation LLMs.
@@ 5.1 Mitigation Strategy based on Quality Quantification
The labeled dataset is randomly split into 80% for training and 20% for testing. Each instance of the labeled dataset is structured as $input=(g_{1},g_{2},\ldots,g_{M},[SEP],e_{1},e_{2},\ldots,e_{N},[SEP]),label=BLEU$, where $g_{i}$ represents tokens in German sentences, $e_{j}$ represents tokens in English translations, and $[SEP]$ denotes the special separation token. Similar to an approach used in Chowdhury et al., (2021), we derive embedding of the entire input sequence from multilingual BERT (with the bert-base-multilingual-uncased pre-trained model), which is then used as input features to eight different supervised learning techniques for BLEU prediction. We train a separate BLEU prediction model for each of the three LLMs, and the testing performance of these BLEU prediction models are summarized in Appendix C. We find that the Bayesian Ridge technique exhibits relatively superior BLEU prediction performance (achieving lower MSE and MAE values).
@@
$\text{Translation Entropy}=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v\in V}p(y_{t}^{v})\log{p(y_{t}^{v})})$ (3)
@@
Finally, we also apply the quality-based mitigation strategy on the Q&A task. For a given question, we analogously derive an entropy score for a model-generated answer to reflect the uncertainty of the probability distribution over all possible answers. For Q&A task, the transformer model generates a candidate answer by predicting the positions of a start token and an end token in a given passage that potentially contains the answer (Rajpurkar et al.,, 2016). The start / end tokens then jointly determine the answer text. Formally, the probability score of a candidate answer is defined as the product of the probabilities associated with the start and end tokens after a softmax transformation, and the entropy score can be computed as
Answer $\text{Answer Entropy}=-\sum_{a\in A}p(y_{a})\log\left\{p(y_{a})\right\}$ (4)
where $a\in A$ is a candidate answer and $p(y_{a})=\frac{p(y_{a}^{start})p(y_{a}^{end})}{\sum_{a\in A}p(y_{a}^{start})p(y_{a}^{end})}$. Same as before, a lower entropy corresponds to a more confident model-generated answer.
We then follow the same settings in Section 3.3 to conduct regurgitative training on both low-performance and high-performance baseline models, taking into account answer entropy. Recall that each article contains multiple training instances (question-answer pairs). Within an article, we rank answers by their entropy scores, from low to high. Across articles, we prioritize articles based on the lowest entropy value among their constituent answers. The results are shown in Figure 10.
@@
We again observe some performance gains of regurgitative training with quality-ranked data, but the improvements tend to be fairly small and dissipate after a sufficient amount of data are added. We repeat this set of experiments but instead prioritize articles based on the average entropy among their constituent answers, and the results remain qualitatively unchanged (Appendix D). To summarize, while the experiments in this section generally support the potential of the quality-based mitigation, we note that the performance gap from training with real data remains substantial. Put differently, the value of real data in training language models cannot be substituted by quality-aware regurgitative training.